### v2

# LLM-based multi-agent poetry generation in non-cooperative environments

Abstract
Despite substantial progress of large language models (LLMs) for automatic poetry generation, the generated poetry lacks diversity while the training process differs greatly from human learning. Under the rationale that the learning process of the poetry generation systems should be more human-like and their output more diverse and novel, we introduce a framework based on social learning where we emphasize non-cooperative interactions besides cooperative interactions to encourage diversity. Our experiments are the first attempt at LLM-based multi-agent systems in non-cooperative environments for poetry generation employing both training-based agents (GPT-2) and prompting-based agents (GPT-3 and GPT-4). Our evaluation based on 96k generated poems shows that our framework benefits the poetry generation process for training-based agents resulting in 1) a 3.0-3.7 percentage point (pp) increase in diversity and a 5.6-11.3 pp increase in novelty according to distinct and novel n-grams. The generated poetry from training-based agents also exhibits group divergence in terms of lexicons, styles and semantics. prompting-based agents in our framework also benefit from non-cooperative environments and a more diverse ensemble of models with non-homogeneous agents has the potential to further enhance diversity, with an increase of 7.0-17.5 pp according to our experiments. However, prompting-based agents show a decrease in lexical diversity over time and do not exhibit the group-based divergence intended in the social network. Our paper argues for a paradigm shift in creative tasks such as automatic poetry generation to include social learning processes (via LLM-based agent modeling) similar to human interaction.
Keywords poetry generation  $\cdot$ social learning  $\cdot$ multi-agent system
1 Introduction
Autonomous agents driven by large language models (LLMs) have made substantial progress in various domains including complex task-solving Li et al. (2024), reasoning Lin et al. (2024); Du et al. (2023), and simulation Wang et al. (2023a). Studies have shown that interactive communication via mono- or multi-agent systems can yield emergent behaviors Park et al. (2023), enhanced task performance Zhuge et al. (2023), better evaluation Chan et al. (2023), and assistance in open-end generation tasks Zhu et al. (2023), to name a few. Despite these advancements, the exploration of creative tasks such as poetry generation utilizing LLM-based agents is still limited Chakrabarty et al. (2023). This paper presents the first experiment on LLM-based multi-agent poetry generation. [^1] The code, data and generated outputs are publicly available at https://github.com/zhangr2021/Multiagent_poetry.git We introduce a framework that emphasizes non-cooperative environments to enhance diversity and novelty in generated poetry both in aggregative mean over time and dynamically across iterations.
Figure 1: Illustration of the predefined social network ($M=4$) and the high-level description of the learning process for training-based agents (GPT-2) and prompting-based agents (GPT-3.5 and GPT-4). The green and red lines in the social network indicate cooperative (+) and non-cooperative (-) interaction between agents respectively.
Why poetry generation?
“Poetry is the rhythmical creation of beauty in words.”
                                                                     – Edgar Allan Poe
Despite advancements in LLMs, generating poetry remains a difficult task due to the complex interplay of style, meaning, and human emotion Chakrabarty et al. (2021); Mahbub et al. (2023). Models need to excel not only in linguistic competence such as understanding semantics and grammar but also in capturing stylistic elements such as rhyme, meter, imagery, and artistic flair to produce human-like poetry Zhipeng et al. (2019); Belouadi and Eger (2023); Ma et al. (2023). Current challenges in poetry generation include: 1) limitations of finetuned models or pipelines that are tailored for specific styles or topics Lau et al. (2018); Van de Cruys (2020); Tian and Peng (2022); 2) LLMs struggling to create diverse and aesthetically pleasing poetry in zero-shot and few-shot scenarios compared to human compositions Sawicki et al. (2023a, b). The complexity and constraints in poetry generation make it a suitable scenario for multi-agent settings, given the known capacity of multi-agent systems to solve complex tasks and the potential to enhance poetry diversity by leveraging “mixtures of multiple poets” Yi et al. (2020).
Why multi-agent system and social learning?
The current process of generating poetry in machine learning and NLP differs substantially from the way in which humans learn and compose poetry. Different from the paradigm of training one single poetry generation model that learns from particular datasets, human beings learn within a social context where they interact with others through communication, be it spoken or written, formal or non-formal Jarvis (2012). This raises the question of whether a more human-like generation approach could improve poetry production. Multi-agent systems, with rich applications in social simulation Park et al. (2023); Chuang et al. (2023), are suitable for our goal where human behavior, such as poetry composition, can be effectively replicated through LLM-based agents that can be situated well in social networks. Furthermore, as one of the most fundamental elements for creative composition, "divergent" thinking ability (to deviate from established norms) during the learning phase is crucial for both human and computational creativity Elgammal et al. (2017); Brinkmann et al. (2023); Wingström et al. (2023). This serves as a rationale for us to adopt a social learning process that supports not only collaboration but also opposition as the main source of deviation in the process Eger (2016); Shi et al. (2019). Moreover, unlike previous studies that focus on weak forms of opposition, such as debates and arguments, which emphasize refining differences through thought correction Chan et al. (2023), our proposed social learning process aims to amplify divergence to enhance diversity.
Why non-cooperative environments?
One aspect of human behavior is non-cooperative interactions in various contexts. For example, the political arena is often characterized by forms of opposition between individual parties or, in a wider context, ‘counter-cultures’ rebelling against the establishment. Moreover, one defining property of literature / art / philosophy is to distinguish oneself from previous or contemporary ‘competitors’. To name some examples, Hemingway’s ‘iceberg’ writing style differs substantially from his predecessors, characterized by a more sentimental writing style Baker (1972); Impressionist artists have challenged the standards of paintings set by the conventional art community with new contents and styles; philosopher Arthur Schopenhauer has strongly argued against the philosophy of Hegel and philosophers often form opposing groups, e.g., ‘Kantians’, ‘Neoplatinists’, etc Janaway (2002). In computational social science, the terms ‘antagonistic’, ‘non-cooperative’ or ‘negative relationships’ are used to describe such behavior Amirkhani and Barshooi (2022). In ML or NLP, such behavior remains rarely explored or leveraged in modeling Gautier et al. (2022); Lei et al. (2022).
How do we utilize LLM-based agents for poetry generation?
We build our social learning framework upon a predefined social network that governs the interactions among agents shown in Figure 1. This network facilitates not only cooperative interactions (“Poets appreciate each other’s work and learn from the others”) but also non-cooperative interactions (“Poets dislike each other’s work and deviate from each other”). We present the learning frameworks based on two different LLMs: 1) prompting-based conversational agents (GPT-4) to answer can our framework benefit in zero-shot or few-shot settings to produce diverse poetry? and 2) training-based agents (GPT-2), where we investigate various training and decoding configurations such as training losses and the number of interactive agents to determine which strategy is the most effective in non-cooperative environments for poetry generation? Our framework comprises three main components: (1) the social network, (2) the learning process and (3) the learning strategy. Our main contribution lies in the development of a novel learning framework that integrates existing methods to extend their scope and effectiveness.
Training-based agents in non-cooperative environments generate poetry of increasing diversity and novelty over time: Our evaluation in Section 5 suggests that our framework benefits the generation process resulting in increasing diversity and novelty according to distinct and novel n-grams for training-based agents. The generated poetry from training-based agents also exhibits group divergence in terms of lexicons, styles and semantics per the predefined group affiliation. Analysis from Section 6 also indicates that non-cooperative conditions boost diversity for prompting-based agents and the potential benefits of employing a more diverse ensemble of models with non-homogeneous agents. But prompting-based agents in our framework do not exhibit group-based divergence of any kind. Moreover, prompting-based agents are prone to generating poetry of homogeneous styles over time.
2 Related work
Our research connects to 1) the interaction of LLM-based agents where we focus on the forms of interaction; 2) the corresponding methods to model interactions, i.e., language model ensemble and controlled text generation (CTG) where CTG techniques are required for our use case in 3) poetry generation.
The interaction of LLM-based agents The form of interactions among agents can be broadly categorized as cooperative and non-cooperative. Very often, agents communicate cooperatively where the aim is to make joint decisions through collaboration such as back-and-forth communication Li et al. (2024), majority voting Hamilton (2023) or a combination of both Zhuge et al. (2023). Non-cooperative interactions, though less prevalent in comparison, can enhance the quality of responses through debates or arguments among agents Chan et al. (2023); Du et al. (2023). Moreover, while most studies focus on static interaction among agents, some researchers delve into the dynamics of agents’ interactions. Liu et al. (2023) propose a dynamic interaction architecture where they utilize an optimization algorithm to select the best agents at inference time. Autogen Wu et al. (2023) also enables dynamic group chat to guide the flow of interaction among agents during ongoing conversations. Others focus on the output dynamics during the interaction, instead. Chuang et al. (2023) find that LLMs tend to align with factual information regardless of their personas and initial states which limits the simulation of opinion dynamics using LLM-based agents. More recently, the dynamics of group interaction is also studied to show that incorporating chain-of-thought reasoning, detailed personas, and finetuning LLMs can enhance agents’ ability to replicate human-like group dynamics Chuang et al. (2024). Our work differs in that 1) we build a social network with group affiliations to obtain a more human-like learning process; 2) we propose a framework that involves two forms of interaction, especially with an emphasis on non-cooperative communication ; 3) we focus on the output dynamics of the generation process under finetuning for training-based agents and consecutive prompting using detailed personas for prompting-based agents.
Language model ensemble and controlled text generation The modeling of our framework requires 1) an ensemble of multiple language models (LMs) and 2) the generated outputs to capture general poetic styles concerning the use case of poetry generation. This involves the areas of LM ensemble and controlled text generation (CTG).
LM ensemble can be divided into 1) conversational ensemble which does not involve parameter training Wang et al. (2022a) and 2) finetuning-based ensemble i.e., neural network ensemble Shazeer et al. (2016) and output ensemble Dekoninck et al. (2023); Jiang et al. (2023). Prompting-based conversational ensemble is often utilized for reasoning tasks where LLMs ensemble their own responses (i.e., self-ensemble) Wang et al. (2022a); Fu et al. (2022). Very recently, Lu et al. (2024) combine multiple small conversational models in a parameter-efficient and interpretable way that outperforms ChatGPT according to their A/B test. Finetuning-based ensembles can operate at the neural network or output level. While neural network ensemble typically requires massive training or finetuning through extensive datasets and resources Shazeer et al. (2016); Jiang et al. (2024), the mixture of smaller modules such as adapters becomes a viable solution in resource-restricted situations Wang et al. (2022b); Chronopoulou et al. (2023).
CTG, the task of generating texts subject to attributes such as emotion Firdaus et al. (2022); Ruan and Ling (2021), topic Dathathri et al. (2019); Wang et al. (2019), toxicity avoidance Liu et al. (2021), style Belouadi and Eger (2023); Shao et al. (2021b), debiasing Dinan et al. (2020); Sheng et al. (2020), etc., is also relevant to our study. CTG can operate at the training/finetuning and inference stage. Similarly to LM ensemble, finetuning with additional modules such as task-related adapters Ribeiro et al. (2021); Lin et al. (2021) is also utilized to gain parameter-efficient controllability. Moreover, for CTG applications to reduce the probability of generating undesirable attributes, in addition to the standard Cross-Entropy loss utilized for text generation finetuning, other loss functions are also explored. Qian et al. (2022) additionally include Contrastive loss which is crucial for the detoxification task but only partially improves the performance of the sentiment control task. Zheng et al. (2023) also employ a Contrastive loss on sequence likelihood to decrease the generation probability of negative samples. In comparison, operation at the inference stage is more viable in the era of LLMs Jiang et al. (2023); Wang et al. (2023b); Dekoninck et al. (2023). Reranking the outputs is a popular solution. For example, Jiang et al. (2023) first rerank the complete candidate outputs from multiple LLMs and then fuse the top-K answers. Reranking the original next-token distribution during the decoding stage is also widely explored such as utilizing discriminators Dathathri et al. (2019) or combining opinions (i.e., output logits) from (anti-)expert models Liu et al. (2021); Dekoninck et al. (2023). Operation during inference, especially at the decoding stage, offers strong controllability over the generated texts with less requirement on time and computational resources. However, it may cause a slight decrease in text quality Dathathri et al. (2019). On the other hand, operations at the training/finetuning stage preserve high-quality text with weaker controllability Zhang et al. (2023).
For our use case of poetry generation, we consider both prompting- and finetuning-based ensemble methods. For the finetuning-based method, we operate jointly 1) at the training stage experimenting with standard Cross-Entropy loss and (or) Contrastive loss to finetune with adapters for better quality texts and efficiency and 2) at the decoding stages to obtain better controllability in the non-cooperative environments.
Automatic poetry generation Early attempts to automatic poetry generation mainly rely on grammatical rules Oliveira (2012), statistical rules Jiang and Zhou (2008); Greene et al. (2010) and neural networks such RNNs Zhang and Lapata (2014); Ghazvininejad et al. (2017); Wöckener et al. (2021), especially RNN-based encoder-decoder architecture Wang et al. (2016); Lau et al. (2018); Yan (2016). More recent models focus on transformer-based architectures Tian et al. (2021); Shao et al. (2021a). Although variants of GPT models have demonstrated outstanding performance in many NLP tasks, mixed opinions are observed on their ability to generate poetry. Studies such as Bena and Kalita (2019); Liao et al. (2019); LC (2022) finetune GPT-2 with additional components such as emotion, form and theme and result in moderate to high-quality poems according to human evaluation. Köbis and Mossink (2021) claim that zero-shot GPT-2 can generate human-like poems where the best poem according to human selection can match human-written ones but without human preselection, machine-generated poems are easily identifiable. Wöckener et al. (2021) point out that similar to RNN-based models, GPT-2 faces problems learning poetry-specific characteristics such as rhyme and meter. To counter such deficiencies, Belouadi and Eger (2023) propose ByGPT5, an end-to-end token-free model conditioned on rhyme, meter and alliteration. The model can outperform larger models such as GPT-2, ByT5, and ChatGPT (GPT3-3.5) according to both automatic and human evaluation. They also construct a customized corpus QuaTrain consisting of large-scale machine-labeled pseudo-quatrains to enlarge the finetuning dataset size. Moreover, Sawicki et al. (2023b) claim that GPT-3 finetuned on 300 poems can successfully generate high-quality poems in a specific author’s style but GPT-3.5 without finetuning leads to undesirable poems written with similar styles. The findings from Sawicki et al. (2023a) also point out that GPT-3.5 and GPT-4, without finetuning, fail to generate poetry in desired styles. Recently, interactive poetry generation has attracted the attention of researchers as it facilitates human-machine collaboration to generate poetry of more diverse styles and better quality under specific constraints Zhipeng et al. (2019); Uthus et al. (2022). Ma et al. (2023) propose a post-polishing system that fine-grains GPT-2 generation based on constraints from humans. CoPoet from Chakrabarty et al. (2022) finetunes pretrained T5 with <instruction, generation> pairs to enable poetry generation according to human instructions. Their study shows that finetuned T5 model is not only competitive to the larger InstructGPT model but also successfully collaborates with humans to produce better poems. In our study, we utilize GPT-2 as our base model as it is a good trade-off between parameter efficiency and language proficiency. Contrary to most poetry generation objectives that optimize few poetry-specific characteristics, we leverage poems with arbitrary styles where we initialize our models with randomly drawn samples from QuaTrain corpus that contain pseudo poetic features. Additionally, we also explore the potential of GPT-3.5 and GPT-4 in an interactive environment enabled by a multi-agent system.
3 Social learning framework for poetry generation
This section introduces our social learning framework for poetry generation. The recent development of LLMs has incentivized various attempts to simulate the social learning processes of human individuals via LLM-based agents Li et al. (2023); Chuang et al. (2023); Gao et al. (2023). Our framework, inspired by such processes, applies a social network-based approach to poetry generation. We investigate whether a more human-like learning process (i.e., social learning) can facilitate poetry generation. We differ from the previous studies in two aspects: 1) We base our architecture on a signed network where agents not only interact in cooperative but also in non-cooperative manners; 2) we introduce the learning framework for prompting-based agents (GPT-3.5 and GPT-4) and training-based agents (GPT-2). Our framework consists of three parts: (1) the social network, (2) the learning process and (3) the learning strategy. We describe these below.
3.1 The social network
We consider a signed social network with $M$ LLM-empowered agents where each link between two agents is associated with a positive or negative sign Leskovec et al. (2010); Eger (2016); Shi et al. (2019). We divide the $M$ agents into two groups as shown in Figure 1. Agents within the same group are referred to as ‘in-group’ agents while agents from different groups are termed ‘out-group’ agents. We expect two types of interaction between agents based on group affiliation: 1) ‘in-group’ agents cooperate closely with one another as ‘friends’ (positive sign); 2) ‘out-group’ agents are ‘foes’ and they adjust their ‘opinions’ in a non-cooperative manner (negative sign). We call the learning process associated with ‘in-group’ agents positive learning and ‘out-group’ associated learning is negative learning. Simultaneous learning from both ‘in-group’ and ‘out-group’ is called joint learning by us.
In our application for poetry generation, the agents are pretrained LLMs. We view the LLMs as different poets belonging to two groups. The ‘in-group’ poets appreciate each other’s work and aim to learn from their styles. Conversely, the ‘out-group’ poets dislike each other’s work and aim to differentiate their works.
3.2 The learning process
Variable Definition
$a_{1},a_{2},\ldots$ agents belong to group A
$b_{1},b_{2},\ldots$ agents belong to group B
M the total number of agents
$\mathcal{A}_{i}$ the target agent $\mathcal{A}_{i}\in\{a_{1},b_{1},a_{2},b_{2},\ldots\}$ where $i\in\{1,2,\ldots,M\}$
$\mathcal{A}_{i},\mathcal{A}^{+},\mathcal{A}^{-}$ the agent tuple: (the target agent, the ‘in-group’ agents of the target agent, the ‘out-group’ agents). E.g., ($a_{1},[a_{2},a_{3}],[b_{1},b_{2}]$)
$P_{{\mathcal{A}_{i}}},P_{*^{+}},P_{*^{-}}$ the conditional probability distribution for the next token of agent $\mathcal{A}_{i}$, agents $*^{+}\in\mathcal{A}^{+}$ and agents $*^{-}\in\mathcal{A}^{-}$
N the total number of generated poems per iteration per agent
T the total number of iterations
t the iteration number $t\in\{1,2,\ldots,T\}$
$o^{\mathcal{A}_{i}}$ a generated poem by agent $\mathcal{A}_{i}$
$O_{t}^{\mathcal{A}_{i}}$ the set of poems generated by agent $\mathcal{A}_{i}$ at iteration t
$S_{t}$ the set of all generated poems at iteration $t$
$F_{\text{generate}}$ a generation function of agent tuple $(\mathcal{A}_{i},\mathcal{A}^{-},\mathcal{A}^{+})$
$F_{\text{update}}$ an updating function based on the latest generated outputs $S_{t}$ and $S_{t-1}$
$t_{g}$ the generation time at the decoding stage
$x_{t_{g}}$ the token at generation time $t_{g}$
$\boldsymbol{x_{t_{g}}}$ the input sequence at generation time $t_{g}$
#$\mathcal{A}$ the number of interactive agents at the decoding stage
Table 1: Notations
for $t\leftarrow 1$ to $T$ do
       Initialize an empty set $S_{t}$ to store the generated poems at iteration $t$;
       foreach agent $\mathcal{A}_{i}$ // $i\in\{1,2,\ldots,M\}$ do
             $O_{t}^{\mathcal{A}_{i}}\leftarrow F_{\text{generate}}(\mathcal{A}_{i},\mathcal{A}^{-},\mathcal{A}^{+})$ to generate $N$ poems;
             Add $O_{t}^{\mathcal{A}_{i}}$ to $S_{t}$;
       end foreach
      foreach agent $\mathcal{A}_{i}$ // $i\in\{1,2,\ldots,M\}$ do
             if $t>1$ then
                   $\mathcal{A}_{i}\leftarrow F_{\text{update}}(S_{t},S_{t-1})$;
            else
                   $\mathcal{A}_{i}\leftarrow F_{\text{update}}(S_{t})$;
             end if
       end foreach
end for
Algorithm 1 Social learning process for poetry generation
Now, we describe the learning process among agents based on the social network shown in Figure 1 and Algorithm 1. We summarize all the notations mentioned below in Table 1. The learning process represents the high-level communication procedure among agents. We begin with pretrained LLM-based agents $a_{1},a_{2},\ldots$ belonging to group $A$ and $b_{1},b_{2},\ldots$ belonging to group $B$ (see Section 4 for details of agent initialization). We further divide the learning process into two phases: the Update phase and the Generate phase, defining the learning strategy $(F_{\text{update}},F_{\text{generate}})$. The two functions jointly organize the positive, negative and joint learning process. $F_{\text{generate}}$ is a generation function that outputs poems $O$. $F_{\text{update}}$ is a learning function that equips agents with the latest knowledge based on the generated poems. We denote an agent as $\mathcal{A}_{i}$ where $i\in\{1,2,\ldots,M\}$. The ‘out-group’ agents for $\mathcal{A}_{i}$ are referred to as $\mathcal{A}^{-}$ and the ‘in-group’ agents as $\mathcal{A}^{+}$. The generation function is thus $F_{\text{generate}}(\mathcal{A}_{i},\mathcal{A}^{-},\mathcal{A}^{+})$.
At each iteration $t$, the learning process starts with the Generate phase. Each agent $\mathcal{A}_{i}$ generates a set of $N$ poems through function $F_{\text{generate}}(\mathcal{A}_{i},\mathcal{A}^{-},\mathcal{A}^{+})$. We iterate over all agents and collect a set of poems $S_{t}$ which consists of all generated poems $O_{t}$ from the current iteration $t$. Then, we let agent $\mathcal{A}_{i}$ update their knowledge and learn cooperatively, non-cooperatively or jointly from each other based on poems from the current iteration $S_{t}$ and the previous iteration $S_{t-1}$. [^2] We utilize the generation from the current and previous iteration to expand the dataset size of the finetuning step where the agent can update to the latest knowledge and prevent potential catastrophic forgetting Biesialska et al. (2020). We denote the update function as $F_{\text{update}}(S_{t},S_{t-1})$. We iterate over all agents $\mathcal{A}_{i}$ for $i\in\{1,2,\ldots,M\}$ until all agents have been updated. We conduct the learning process $T$ times.
3.3 The learning strategy
As mentioned in Section 3.2, the learning process involves positive, negative and joint learning which operates at the Generate and the Update phase. We now discuss the detailed learning strategy for training-based agents and prompting-based agents. For training-based agents, the learning strategy contains finetuning strategies for the Update phase & decoding strategies for the Generate phase. For prompting-based agents, both phases are conducted via prompting. Table 2 summarizes the learning strategies for both types of agents.
Type of agents Strategy Positive learning Negative learning Joint learning
($\mathcal{A}_{i},\mathcal{A}^{+}$) ($\mathcal{A}_{i},\mathcal{A}^{-}$) ($\mathcal{A}_{i},\mathcal{A}^{+},\mathcal{A}^{-}$)
training-based decoding - $P_{{\mathcal{A}_{i}}}$, $P_{*^{-}}$ $P_{{\mathcal{A}_{i}}},P_{*^{+}},P_{*^{-}}$
finetuning $\mathcal{L}_{\text{CE}}$ 1) $\mathcal{L}_{\text{CL}}$
2) conditioned $\mathcal{L}_{\text{CE}}$
prompting-based prompting chain-prompting joint-prompting
Table 2: Learning strategies for training-based agents and prompting-based agents. $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{CL}}$ represent Cross-Entropy loss and Contrastive loss. $\mathcal{A}_{i},\mathcal{A}^{+},\mathcal{A}^{-}$ denotes the target agent, the ‘in-group’ agent and the ‘out-group’ agent of the target agent $\mathcal{A}_{i}$. $P_{{\mathcal{A}_{i}}},P_{*^{+}},P_{*^{-}}$ are the conditional probability distribution for the next token of agent $\mathcal{A}_{i}$, agents $*^{+}\in\mathcal{A}^{+}$ and agents $*^{-}\in\mathcal{A}^{-}$ defined in Equation (1).
3.3.1 Training-based agents
We first introduce the decoding strategy for the Generate phase where we utilize reranking techniques. We then detail the finetuning strategy for the Update phase where we employ the Contrastive loss and the standard Cross-Entropy loss.
Decoding strategy at the Generate phase
[…]
Given the input sequence at generation time $t_{g}$ ($g$ indicates the generation stage), denoted as $\boldsymbol{x_{<t_{g}}}$, we predict the next token $x_{t_{g}}$ generated by the target agent $\mathcal{A}_{i}$ through combining the outputs from $\mathcal{A}^{+}_{\#}$ and $\mathcal{A}^{-}_{\#}$. We first obtain the conditional logit scores of all models denoted by $l_{\mathcal{A}_{i}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}}),l_{*^{+}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}}),l_{*^{-}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}})$, where $*^{+}\in\mathcal{A}^{+}_{\#}$ is an agent belonging to the interactive ‘in-group’ agents $\mathcal{A}^{+}_{\#}$ and $*^{-}\in\mathcal{A}^{-}_{\#}$; $l_{*}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}})\in\mathbb{R}^{|\mathcal{V}|}$ and $\mathcal{V}$ is the vocabulary. The probability distribution of the next token over the vocabulary $\mathcal{V}$ is $P_{*}({x_{t_{g}}|\boldsymbol{x_{<t_{g}}}})=\text{softmax}[l_{*}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}})]$. The probability distribution of the next token $\hat{P}_{\mathcal{A}_{i}}({x_{t_{g}}|\boldsymbol{x_{<t_{g}}}})$ is thus given by
$\begin{split}\hat{P}_{\mathcal{A}_{i}}({x_{t_{g}}|\boldsymbol{x_{<t_{g}}}})=&\text{softmax}\{l_{\mathcal{A}_{i}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}})+\\ &\alpha[\frac{\sum_{\mathcal{A}^{+}_{\#}}{l_{*^{+}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}})}}{\rvert\mathcal{A}^{+}_{\#}\rvert}-\frac{\sum_{\mathcal{A}^{-}_{\#}}{l_{*^{-}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}})}}{\rvert\mathcal{A}^{-}_{\#}\rvert}]\}\end{split}$ (1)
The next token $x_{t_{g}}$ is assigned with high probability if the probability is high under both $P_{{\mathcal{A}_{i}}}$ and $P_{*^{+}}$ and low under $P_{*^{-}}$. Moreover, if we replace $P_{*^{+}}$ with $P_{{\mathcal{A}_{i}}}$, the process considers ‘out-group’ agents $\mathcal{A}^{-}$ only which solely models negative learning. This decoding strategy can model both negative learning ($P_{{\mathcal{A}_{i}}},P_{*^{-}}$) and joint learning ($P_{{\mathcal{A}_{i}}},P_{*^{-}},P_{*^{+}}$) as shown in Table 2.
Finetuning strategy at the Update phase We discuss the finetuning strategies based on the learning relationships, i.e., positive and joint learning as shown in Table 2.
  • Positive learning: we utilize the conventional finetuning method with Cross-Entropy loss ($\mathcal{L}_{\text{CE}}$) to finetune agent $\mathcal{A}_{i}$ with poems $o\in O_{\mathcal{A}_{i}}\cup O_{\mathcal{A}^{+}}$ (poems generated from $\mathcal{A}_{i}$ and $\mathcal{A}^{+}$). The loss function for the $j^{th}$ poem $o_{j}$ in a mini-batch is thus
    $\mathcal{L}_{\text{CE}}(\mathcal{A}_{i},o_{j})=-\sum_{t_{g}=1}^{\mathcal{T}}\log(P_{\mathcal{A}_{i}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}}))$ (2)
    where $\mathcal{T}$ denotes the number of tokens for the poem $o_{j}$. $P_{\mathcal{A}_{i}}(x_{t_{g}}|\boldsymbol{x_{<t_{g}}})$ is the conditional distribution of the token at time $t_{g}$ for poem $o_{j}$ given the previous sequence $\boldsymbol{x_{<t_{g}}}$.
  • Joint learning with Contrastive loss: Our social network design well suits the setting for Contrastive learning where poems from ‘in-group’ agents are positive samples and poems from ‘out-group’ agents are negative samples. We adopt Contrastive learning to pull closer the semantic representation of ‘in-group’ samples and push apart that of ‘out-group’ samples. We implement the Contrastive loss SimCSE proposed by Gao et al. (2021). For a mini-batch containing samples from $\mathcal{A}_{i},\mathcal{A}^{+},\mathcal{A}^{-}$, let $(o_{j}^{\mathcal{A}_{i}},o_{j}^{\mathcal{A}^{+}},o_{j}^{\mathcal{A}^{-}})$ denote the $j^{th}$ paired triple and $(\boldsymbol{h_{j}},\boldsymbol{h_{j}}^{+},\boldsymbol{h_{j}}^{-})$ be its representation. The Contrastive loss is thus
    $\mathcal{L}_{\text{CL}}(\mathcal{A}_{i},(\boldsymbol{h_{j}},\boldsymbol{h_{j}}^{+},\boldsymbol{h_{j}}^{-}))=-\log\frac{e^{\text{sim}(h_{j},h_{j}^{+})/\tau}}{\sum_{k=1}^{Q}\left(e^{\text{sim}(h_{j},h_{k}^{+})/\tau}+e^{\text{sim}(h_{j},h_{k}^{-})/\tau}\right)}$ (3)
    where $Q$ is the size of mini-batch, $\tau$ is the temperature and $sim(h_{1},h_{2})$ is the cosine similarity $\frac{h_{1}^{\mathsf{T}}h_{2}}{\|h_{1}\|\cdot\|h_{2}\|}$. Here, we experiment with Contrastive loss (a) alone and (b) jointly with $\mathcal{L}_{\text{CE}}$ for ‘in-group’ poems.
  • Joint learning with conditioned Cross-Entropy loss: We utilize the style constraints Belouadi and Eger (2023) using conditions <positive> and <negative> for poems generated by ‘in-group’ agents and ‘out-group’ agents. We then finetune the agent $\mathcal{A}_{i}$ with Cross-Entropy loss.
3.3.2 Prompting-based agents
The learning strategy for prompting-based agents relies on prompting. Our prompts are constructed with three modules: 1) a profile module, which defines the role of $\mathcal{A}_{i}$; 2) a memory module, which stores the generated poems; and 3) an action module, which completes the generation task. Table 8.1 in the appendix contains the prompts for both prompting strategies. For prompting-based agents, the Update and Generate phases do not function in isolation. $F_{\text{update}}$ updates the profile module during prompting based on the generated poems from previous iterations. Similar to the design of training-based agents, we can update the knowledge of agents based on different learning relationships:
  • The chain-prompting strategy for isolated positive & negative learning: For chain-prompting, we update the knowledge of $\mathcal{A}_{i}$ based on its relationships with other agents in a separate manner. At iteration $t$, we first update the profile of $\mathcal{A}_{i}$ with poems generated from $\mathcal{A}^{+}$ at time $t-1$. The process is denoted as $F_{\text{update}}(\mathcal{A}_{i},\mathcal{A}^{+})$. $\mathcal{A}_{i}$ thus generates a poem $o^{\mathcal{A}_{i}}$ based on the positive learning results (example prompt: “Please read the poems from your friends carefully and compose similarly to your friend.”). We then update the profile of $\mathcal{A}_{i}$ with the poem $o^{\mathcal{A}_{i}}$ and a poem $o^{\mathcal{A}^{-}}$ sampled from the previous iteration $t-1$. This process is denoted as $F_{\text{update}}(\mathcal{A}_{i},\mathcal{A}^{-})$. $\mathcal{A}_{i}$ thus recomposes the poem $o^{\mathcal{A}_{i}}$ based on the negative learning results (example prompt: “Please rewrite your poem to compose dissimilarly to your foe.”).
  • Joint prompting for joint learning: Joint prompting updates the profile with poems generated by $\mathcal{A}^{+}$ and $\mathcal{A}^{-}$ at the same time, denoted as $F_{\text{update}}(\mathcal{A}_{i},\mathcal{A}^{-},\mathcal{A}^{+})$.
4 Experiments
4.1 Agent initialization
Instance Rhyme Meter alliteration
Who hath such beauty seen
In one that changeth so?
Or where one’s love so constant been,
Who ever saw such woe?
ABAB iambus 0.11
Would rather seek occasion to discover
How little pitiful and how much unkind,
They other not so worthy beauties find.
O, I not so! but seek with humble prayer
ABBC iambus 0.05
Of pearl and gold, to bind her hands;
Tell her, if she struggle still,
I have myrtle rods at will,
For to tame, though not to kill.
ABBB iambus 0.10
Table 3: Instances from QuaTrain corpus.
Initialization for training-based agents We choose $M=4$ for our initial experiments. We first pretrain GPT-2 (medium) with a subset of the random QuaTrain corpus of size 123K (nearly 1/6 of QuaTrain corpus). QuaTrain consists of machine-labeled pseudo-quatrains that are consecutive sequences with four lines extracted from real human poems as shown in Table 3. The poems follow poetic characteristics, i.e., rhyme, meter and alliteration. We pre-select the QuaTrain instances excluding those that are semantically similar ($>0.7$) with each other based on pairwise cosine similarity of sentence embeddings calculated using SBERT Reimers and Gurevych (2019). We first pretrain GPT-2 further with 720 training steps using the pre-selected QuaTrain dataset. More details of the pretraining and the loss curve are shown in Section 8.2. We then finetune the pretrained model with four randomly selected subsets of size 7.5k from our 123K subcorpus to obtain different initializations of four agents. The initialization step prepares models with a preliminary understanding of poetic structures and characteristics essential for subsequent learning phases.
Initialization for prompting-based agents For prompting-based agents, we randomly sample QuaTrain instances and initialize the agent with chain-prompting and joint-prompting under the predefined profile shown in Table 12 and Table 13.
4.2 Experimental setup
RQ1 Para_decoding Para_finetuning Description
#$\mathcal{A}$ $\alpha$ $\mathcal{L}_{\text{CE}}$ $\mathcal{L}_{\text{CL}}$ conditioned
#$\mathcal{A}$ during decoding {2,3,4} 2 X The number of agents involved during decoding is varied.
1) #$\mathcal{A}$ = 2: negative decoding + positive finetuning
2) #$\mathcal{A}$ > 2: joint decoding + positive finetuning
$\alpha$ 2 {0, 1, 1.5, 2, 2.5} X The scaling parameter $\alpha$ during decoding is varied.
$\alpha>0$: negative decoding + positive finetuning
$\alpha=0$: positive finetuning only, ‘echo chamber’
finetuning strategy 2 2 X Different training loss applied for $\alpha$ = 2 (with negative decoding).
1) $\mathcal{L}_{\text{CE}}$ alone: positive training with negative decoding
2) $\mathcal{L}_{\text{CL}}$ or conditioned: joint training with negative decoding
X
X
X X
X X
Table 4: Experimental setup for training-based agents. Para_decoding and Para_finetuning represent parameters during the decoding and finetuning stage. #$\mathcal{A}$ is the number of agents. $\alpha$ is the scaling parameter in Equation (1).
For training-based agents, we design experiments to explore how the parameters from the finetuning stage (i.e., loss functions) and the decoding stage (number of agents and the scaling parameter) affect the dynamics of generation. We summarize the detailed setup in Table 4.
For prompting-based agents, we design experiments with different prompting strategies, i.e., chain-prompting and joint-prompting using both GPT-3.5 (gpt-3.5-turbo) and GPT-4 (gpt-4-turbo).
4.3 Evaluation
We first conduct automatic evaluation where we study the generation dynamics of our framework from lexical perspectives. We study lexical novelty and diversity, as novelty and diversity are crucial indicators for creative tasks such as poetry generation. We then study the dynamics from semantic (semantic similarity) perspectives. Lastly, we evaluate the poems qualitatively where we directly compare the generated poetry.
Metric for lexical diversity and novelty. We measure the lexical diversity using the percentage of distinct uni-grams (distinct-1) and bi-grams (distinct-2) following the definition by Su et al. (2022); Tevet and Berant (2021). The formulation is given as: unique $\frac{\text{unique n-grams}(O)}{\text{total n-grams}(O)}$, where $O$ is the set of generated poems to be evaluated. We adopt the measure of novelty (novelty-1 and novelty-2) by McCoy et al. (2023); Shen et al. (2020) where we calculate the number of new uni-/bi-grams that do not stem from the pretraining set and rescale them with the total number of generated tokens. Thus, novelty reflects the lexical difference between the generated poems and the training set, while diversity indicates the token variety among the generated poems.
Metric for group-based semantic similarity. Following our group affiliation, we examine the group dynamics of the agents by their semantics. For any pair of poems sampled from the same iteration $t$, we calculate the semantic similarity of the paired instances by computing the cosine similarity of the embeddings retrieved from SBERT Reimers and Gurevych (2019). We then aggregate the similarity scores per iteration by their group affiliation (i.e., ‘in-group’ and ‘out-group’) defined in Figure 1.
5 Experiment results
5.1 Automatic Evaluation: the generation dynamics of agents
We generate 400 poems using the same set of decoding parameters for training-based agents or the same prompt templates for prompting-based agents from each agent for every iteration. [^4] See Section 8.2 in the appendix for more details on parameter setting. In total, we obtain more than 96k generated poems. We report the lexical diversity (distinct-1 and distinct-2), novelty (novelty-1 and novelty-2), and group-based semantic similarity defined in Section 4.3.
Our main findings are: 1) according to lexical level comparison (i.e., distinct and novel uni-/bi-grams), our framework benefits training-based agents resulting in increasing diversity and a higher level of novelty; 2) according to pairwise semantic similarity averaged per group affiliation, we observe group divergence in semantics for training-based agents, especially ‘out-group’ divergence due to operation at the decoding stage; 3) prompting-based agents generate poems of more diverse lexicons at $t=1$ but they tend to output poems of homogenous styles over time. [^5] Due to resource limit, we do not conduct multiple runs for all experiments. We study the stability of our statistics in Section 6.1. This section is based on one single run of experiments.
5.1.1 Increasing diversity and novelty according to distinct and novel n-grams for training-based agents
We compute distinct-1/2 and novelty-1/2 with all generated poems and average over all agents for every iteration $t$. The results are shown in Table 5 and Figure 2.
RQ1: how do different learning strategies affect the diversity and novelty of training-based agents?
Figure 2: Dynamics of agent diversity and novelty over varying training parameters. The degree of diversity is measured by the percentage of distinct uni-grams (distinct-1) and bi-grams (distinct-2) in the generated poems. The degree of novelty is measured by the number of novel uni-grams (novelty-1) and bi-grams (novelty-2) in the generated poems compared to that in training data scaled by the total number of generated tokens. (a) The effect of scaling parameter $\alpha$ in Equation (1). (b) The effect of the number of interactive agents #$\mathcal{A}$ during the decoding stage. (c) The effect of finetuning strategies: $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{CL}}$ indicate Cross-Entropy loss and Contrastive loss. Prefix refers to the conditioned finetuning. The horizontal red dashed line indicates the state of initial agents at iteration 0.
[…]
varying $\alpha$
0 2 $\mathcal{L}_{\text{CE}}$ 0.120 0.665 0.035 0.139
1 2 $\mathcal{L}_{\text{CE}}$ 0.154 0.705 0.062 0.202
1.5 2 $\mathcal{L}_{\text{CE}}$ 0.164 0.713 0.077 0.222
2 2 $\mathcal{L}_{\text{CE}}$ 0.167 0.713 0.092 0.237
2.5 2 $\mathcal{L}_{\text{CE}}$ 0.154 0.698 0.105 0.234
varying #$\mathcal{A}$
2 2 $\mathcal{L}_{\text{CE}}$ 0.167 0.713 0.092 0.237
2 3 $\mathcal{L}_{\text{CE}}$ 0.123 0.671 0.037 0.158
2 4 $\mathcal{L}_{\text{CE}}$ 0.171 0.714 0.161 0.264
varying training loss
2 2 $\mathcal{L}_{\text{CE}}$ 0.167 0.713 0.092 0.237
2 2 $\mathcal{L}_{\text{CE}}$ + $\mathcal{L}_{\text{CL}}$ 0.165 0.703 0.103 0.231
2 2 $\mathcal{L}_{\text{CL}}$ 0.156 0.753 0.054 0.160
2 2 $\mathcal{L}_{\text{CE}}$ + prefix 0.159 0.696 0.102 0.217
Initialization 0.171 0.785 0.061 0.190
Table 5: Diversity and novelty results in aggregative mean for training-based agents. Distinct-1 and distinct-2 are the percentage of distinct uni-/bi-grams. Novelty-1 and novelty-2 reflect the number of new uni-/bi-grams that do not appear in the training set rescaled by the total number of generated tokens. The highest value in each experimental setting is highlighted in bold. $\alpha$ represents the decoding scaling parameter; #$\mathcal{A}$ is the number of interactive agents at decoding stage; $\mathcal{L}$ represents the loss function during finetuning. Initialization indicates the states of initial agents at iteration 0.
Figure 2 shows the dynamics of agent diversity and novelty under varying training parameters and Table 5 shows the results of different experimental setups for diversity and novelty averaged over all iterations. We observe that:
  • the effect of negative decoding strategy with the scaling parameter $\alpha$.
    • Diversity Negative decoding combined with positive finetuning ($\alpha$ >0, $\mathcal{L}_{\text{CE}}$) strategy leads to increasing diversity over time though the level of diversity is below the initial state at $t=0$. Aggregatively, results from Table 5 (varying $\alpha$) suggest that compared to the case without negative decoding (i.e., $\alpha=0$), negative decoding strategy under varying $\alpha$ ranging from 1 to 2.5 yields 3.4 to 4.7 percentage points (pp) increase in distinct-1 and 3.3 to 4.8 pp increase in diversity measured by distinct-2. Dynamically, the results from Figure 2(a) suggest that the lexical diversity of generated poems with negative decoding depicts an increasing trend from $t=1$ to $t=4$ for all $\alpha>0$ measured by both distinct-1 (with a maximum increase of 3.7 pp) and distinct-2 (with a maximum increase of 3.0 pp) while for $\alpha=0$ (i.e., without negative decoding), both diversity measures decrease slightly. Worth noting is that both distinct-1 and distinct-2 are below the diversity level measured at $t=0$ (shown as the red dashed line in Figure 2(a) and the last row in Table 5), especially distinct-2.
    • Novelty Negative decoding combined with positive finetuning ($\alpha$ >0, $\mathcal{L}_{\text{CE}}$) strategy boosts novelty over time resulting in more novel generation compared to the initial state at $t=0$. The last two columns in Table 5 show that the negative decoding strategy, i.e., $\alpha>0$, can boost novelty in the aggregative mean by a maximum of 7.0 pp in novelty-1 and a 9.8 pp increase in novelty-2 compared to the case without negative decoding, i.e., $\alpha=0$. Dynamically, results from Figure 2(a) suggest a sharper increase over iterations for all $\alpha>0$ measured by both novelty-1 (with a maximum increase of 5.6 pp) and novelty-2 (with a maximum increase of 11.3 pp) compared to the results for $\alpha=0$.
  • the effect of the number of agents (#$\mathcal{A}$) involved at the decoding stage
    • Diversity As shown in Table 5, #$\#\mathcal{A}=4$ yields the highest diversity level according to distinct-1 and distinct-2 with #$\#\mathcal{A}=2$ achieving similar performance. Dynamically, diversity increases over iteration for paired agents (#$\#\mathcal{A}=2$ or $4$) at the decoding stage. However, for #$\#\mathcal{A}=3$, we observe a decreasing trend in diversity with much lower level of diversity compared to the case for #$\#\mathcal{A}=2$ or $4$.
    • Novelty We observe a greater gain for novelty at #$\#\mathcal{A}=4$. This is evident: 1) Table 5 shows 6.9 pp increase in aggregative mean for #$\#\mathcal{A}=4$ compared to #$\#\mathcal{A}=2$; 2) Figure 2(b) indicates a shaper increasing trend at #$\#\mathcal{A}=4$ especially for novelty-1. Both novelty-1 and novelty-2 are above the initial state at iteration 0 which suggests a boost in novelty over all time. However, we observe less novelty for #$\#\mathcal{A}=3$, which is similar to the case for diversity.
  • the effect of finetuning strategy. The decoding parameters #$\mathcal{A}$ and $\alpha$ are fixed and we experiment with varying finetuning losses. As suggested by Figure 2(c), the most effective finetuning strategy according to the dynamics of diversity and novelty is $\mathcal{L}_{\text{CE}}$ (i.e., positive finetuning using the Cross-Entropy loss) which presents an observable upward trend. Finetuning using $\mathcal{L}_{\text{CL}}$ (i.e., joint finetuning using Contrastive loss) yields slightly better diversity according to distinct-2. We also observe minor improvement in novelty for strategy $\mathcal{L}_{\text{CL}}$ $+$ $\mathcal{L}_{\text{CE}}$ (i.e., joint finetuning using both losses) in aggregative mean shown in Table 5. However, dynamically we do not spot any increase over time for both cases. Conditioned finetuning (i.e., Prefix) also fails to bring improvements.
To sum up, our framework can lead to increasing diversity and a higher level of novelty: 1) negative decoding combined with positive finetuning ($\alpha$ >0, $\mathcal{L}_{\text{CE}}$) is the most effective combination of decoding and finetuning strategies; 2) the increase in an even number of agents can improve the results, especially for novelty; 3) in our experiment, positive finetuning (i.e., finetuning using Cross-Entropy loss alone) is more effective overall both in aggregative mean and dynamically compared to other finetuning strategies.
RQ2: how do different prompting strategies affect the diversity of prompting-based agents?
Figure 3: Dynamics of diversity for prompting-based agents over varying prompting strategies based on GPT-3.5 and GPT-4.
model strategy distinct-1 distinct-2
GPT-3.5 chain 0.352 0.771
GPT-3.5 joint 0.321 0.739
GPT-4 chain 0.404 0.876
GPT-4 joint 0.336 0.817
Table 6: Diversity results in aggregative mean for prompting-based agents. Distinct-1 and distinct-2 are the percentage of distinct uni-/bi-grams.
As prompting-based agents do not involve further pretraining, novelty metrics, which involve comparison to the pretraining dataset, are thus undefined. Therefore, we only study the lexical diversity of the generated poetry. Figure 3 shows the dynamics of diversity over varying prompting strategies for agents based on GPT-3.5 and GPT-4.
Do we observe an increasing trend for prompting-based agents similar to that of the training-based agents? Different from the trend we observe for training-based agents, prompting-based agents exhibit a sharp increase from $t=1$ to $t=2$ with a maximum of 6.3 pp increase in distinct-1 and 10 pp in distinct-2 for nearly all experiments. GPT-3.5 under chain-prompting is an exception where we observe a constant decreasing trend in distinct-2. However, the increment in lexical diversity pauses when $t>2$ where we yield slightly decreasing trends for nearly all experiments. GPT-3.5 under joint-prompting is an exceptional case where the increasing trend continues mildly. We examine the effect of positive and negative learning strategies in separation in Section 6.2.
Which prompting strategy and base model perform better according to lexical diversity? Both Figure 3 and Table 6 indicate that GPT-4 under chain-prompting generates the most lexically diverse poetry compared to other settings. In general, the chain-prompting strategy performs better than joint-prompting according to distinct-1 and distinct-2. However, GPT-4 does not always outperform GPT-3.5 as suggested by the aggregative mean in Table 6 where GPT-3.5 under chain-prompting strategy delivers the second best performance according to distinct-1.
For prompting-based agents, our framework only benefits the generation process in a limited manner (when $t=1,2$) according to lexical diversity. Worth noting is that prompting-based agents have an overall higher percentage of unique uni-grams distinct-1 and bi-grams distinct-2 shown in Table 6, especially for distinct-1 with below 20 pp for training-based agents and over 40 pp for prompting-based agents.
5.1.2 Group divergence in semantics
Figure 4: Divergence of training-based agents measured by mean of pairwise semantic similarity over varying training parameters. (a) The effect of scaling parameter $\alpha$ in Equation (1). (b) The effect of the number of interactive agents #$\mathcal{A}$ during the decoding stage. (c) The effect of finetuning strategies: $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{CL}}$ indicate Cross-Entropy loss and contrastive loss. Prefix refers to the conditioned finetuning. The solid line and dashed line represent semantic similarity measured for ‘in-group’ and ‘out-group’ affiliations respectively.
RQ3: how do different learning strategies affect the group dynamics of training-based agents?
  • Observable group dynamics for positive training (Cross-Entropy loss) with negative decoding. Figure 4 shows the mean semantic similarity based on group affiliations for different scaling parameters $\alpha$, the number of agents #$\#\mathcal{A}$ involved during the decoding stage and different finetuning strategies. The solid line represents semantic similarity measured for ‘in-group’ agents and the dashed line for ‘out-group’ agents. Overall, we observe a divergence between ‘in-group’ and ‘out-group’ similarity for Cross-Entropy loss with negative decoding under varying scaling parameters $\alpha$ and different numbers of agents #$\#\mathcal{A}$. The effects of parameters vary: 1) Figure 4(a) exhibits the dynamics for different $\alpha$. We observe a divergence between the semantic similarity of ‘in-group’ and ‘out-group’ where particularly, ‘out-group’ similarity decreases over iterations. $\alpha=0$ represents the case for ‘echo chambers’ where only positive finetuning is considered (i.e., agents only talk to their ‘in-group’). For $\alpha=0$, the agents echo their own ‘thoughts’ resulting in an overall higher level of similarity for both ‘in-group’ and ‘out-group’ compared to $\alpha>0$. For $\alpha>0$, we yield an 8.8 pp decrease in semantic similarity for ‘out-group’ which is 4.7 pp greater in divergence compared to the case for $\alpha=0$ (4.1 pp in total); 2) we observe from Figure 4(b) that interaction involving more agents during the decoding stage has a slightly positive influence on group divergence. #$\#\mathcal{A}=4$ yields a mild increase with 2.2 pp in ‘in-group’ semantic similarity and an 11.0 pp decrease in ‘out-group’ similarity (13.2 divergence in total). In contrast, #$\#\mathcal{A}=2$ results in an increase with 2.8 pp for ‘in-group’ and 9 pp decrease for ‘out-group’ (11.8 divergence in total). Overall, #$\#\mathcal{A}=4$ exhibits a lower level of similarity compared to #$\#\mathcal{A}=2$.
  • Inseparable ‘in-group’ and ‘out-group’ dynamics resulting from other joint finetuning strategies. Figure 4(c) shows the outcome for different finetuning strategies involving multiple losses $\mathcal{L}$ and conditioned finetuning (i.e., Prefix). Except for the case using $\mathcal{L}_{\text{CE}}$ alone as the finetuning loss (i.e., positive finetuning defined in Table 4), all other cases with joint finetuning exhibit inseparable dynamics between ‘in-group’ and ‘out-group’ similarity. We suspect that for contrastive learning ($\mathcal{L}_{\text{CL}}$), a negative pair built purely based on group affiliation fails to provide enough contrastivity considering that we initiate the agents in a random manner. Such random initialization may affect the results for conditioned finetuning as well.
Figure 5: Group divergence of prompting-based agents measured by the mean of pairwise semantic similarity over varying prompting strategies and base model. The solid line and dashed line represent semantic similarity measured for ‘in-group’ and ‘out-group’ affiliations respectively.
RQ4: how do different prompting strategies affect the group dynamics of prompting-based agents?
Undesirable increasing semantic similarity from ‘out-group’ agents Figure 5 shows the group divergence of prompting-based agents measured by the mean pairwise semantic similarity over varying prompting strategies and base models. The solid line represents the semantic similarity measured for ‘in-group’ agents and the dashed line for ‘out-group’ agents. We observe an increasing similarity for both ‘in-group’ and ‘out-group’ agents where the greatest group divergence is observed at $t=1$. Over time, the agents tend to generate semantically similar poetry for both GPT-3 and GPT-4. Moreover, we also notice that prompting-based agents generate poetry of homogeneous styles over time which coincides the finding of Sawicki et al. (2023a). We discuss this point in more detail in Section 5.2.
5.2 Qualitative evaluation
t group A group B
training-based agents ($\alpha=2,\#\mathcal{A}=2$, $\mathcal{L}_{\text{CE}}$)
0 That will be and the world shall give place to one, whom death has never forgotten. Here thy dearest child is alone: A sad birth thou seekest still more mourn Thou, the child of Israel - thou who dost keep the Lord forever; And every day we mourn thy loss.
1 vultures flying; their cries in one another’s midst: And when thou seeest a poor child, Who’d the hand that’s to feed him must? The children of the world have been lulled to sleep. In sorrow’s night their eyes fail to see - no tears are shed, but only dreams; And they will say that nothing is more beautiful
4 Pretend, I know not where: But this is what has made me love thee so; For thou art a child of my breast, The affection I have felt for her in youth’s day. reputations will be lost; lives of men, women and children shall be gone. But the memory of my life on earth Will grow in faithfulness like a tree-top That grows new leaves that come to rest Where its roots once
prompting-based agents (GPT-3.5 chain-prompting)
0 Tiny hands and curious hearts, Youthful souls, creating art, Innocent laughter fills the air, Children’s world, free from care, Growing up, but never old, Their stories waiting to be told. Youth’s fleeting flame burns bright and bold, Innocence wrapped in dreams untold. Carefree laughter, hearts pure and free, Embracing life’s wonder, wild and carelessly. Oh, to be young, to feel so alive, In a world where anything can thrive.
1 Tears, like a river, flowed Father’s decree, "Stay away, my child" Through the wild woods he roamed The wind, his sole confidant He races, longing for home, heart mild Arise, oh heart! Unveil thy pure essence, as we danced in our vibrant youth.
4 In the moon’s gentle light, I dance with shadows in the night, With a heart filled with a melody, I roam through nature’s symphony. Lost in the shadows of a world unknown, Drifting through the silence all alone, Seeking solace in the whispers of the night, Longing for a glimpse of dawn’s soft light.
Table 7: Generated poetry based on group affiliation from training-based and prompting-based agents at different iteration $t$. For prompting-based agents, the baseline at $t=0$ is generated using one simple prompt (‘Please generate a poem about children or youth”). We highlight words containing historical components in italics. Words that rhyme are highlighted in bold. Words that mildly rhyme are colored in gray. Grammatical errors are marked in red.
Table 7 contains examples of generated poetry from training-based agents under positive finetuning and negative decoding strategy (i.e., $\alpha=2,\#\mathcal{A}=2$, $\mathcal{L}_{\text{CE}}$) and from prompting-based agents using GPT-3.5 under the chain-prompting strategy. We select examples composed with similar themes, i.e., child or youth. Poems generated at $t=0$ are considered the baselines for both training-based and prompting-based frameworks.
For training-based agents, at $t=0$ the generated poems often contain historical spellings (e.g., ‘thy’ and ‘thou’) and historical morphology terms (e.g., ‘seekest’ and ‘dost’). Apart from the semantic divergence discussed in the previous section, we observe a divergence in word choices over time. For example, we study the poems generated by training-based agents with settings $\alpha=2,\#\mathcal{A}=2$, $\mathcal{L}_{\text{CE}}$ where we calculate the percentage of poems that contain historical spellings and historical morphology terms. We find that over 26% of poems generated by agents in group A contain historical spellings or morphology terms compared to only 10% of poems by agents from group B. Moreover, the percentage of poems with historical languages for group A is stable at a level of 26% over iterations while for group B, the percentage steadily decreases over time by nearly 6 pp. The word frequency of poems from group A and group B also suggests such divergence. For example, words such as “mind, thy, thee, nature, art, power, happy, hath, young, pleasant, friend, …” are more frequent in group A and less frequent in group B while in group B the more frequent words are “god, lord, light, sun, sky, sea, land, soul, children, dream, …”. For prompting-based agents, in contrast, we observe more diverse lexicons for both groups compared to training-based agents but the two groups in prompting-based setting hardly diverge in terms of vocabulary or topics when $t>1$. This is also suggested by Figure 5. Moreover, prompting-based agents tend to generate poems of homogenous styles over time. As shown in Table 7, poems generated from prompting-based agents excessively focus on rhymes which makes the generated poetry merely superficially human-like. This also suggests a poor understanding of poetry for GPT-3.5 and GPT-4 in a zero-shot setting. Even though GPT-3.5 and GPT-4 can adopt historical texts well Zhang et al. (2024), they never pick up the historical expressions from the initial poetry as the training-based agents do. Apart from the ‘obsession’ for rhyming, GPT-3.5 and GPT-4 also tend to generate poems using similar beginning phrases such as “Beneath/Under XXX, In the XXX, Lost in XXX”, especially when $t>1$. The generated poems from GPT-3.5 and GPT-4 contain fewer grammatical errors than training-based agents, though training-based agents generate poems of more diverse styles and topics in comparison.
6 Discussion and analysis
6.1 How stable are the simulation results?
Due to resource constraints, we do not execute multiple simulations for all experiment settings. Instead, we study the stability of our experiments using two experiment settings for training-based agents, i.e., $\alpha=0$ and $\alpha=2$, and one experiment setting for prompting-based agents, i.e., GPT-3,5 chain-prompting. We rerun the experiments three times under the same parameters (or prompt templates for prompting-based agents) and initialization. We then yield three sets of statistics and calculate the standard deviation as our stability measure. We study the stability from two perspectives: 1) stability of the aggregative mean and 2) dynamic stability.
model & setting distinct-1 distinct-2 novelty-1 novelty-2
training-based $\alpha=0$ 0.001 (.120) 0.002 (.664) 0.001 (.034) 0.003 (.136)
$\alpha=2$ 0.003 (.164) 0.004 (.709) 0.004 (.095) 0.005 (.238)
prompting-based GPT-3.5 chain-prompting 0.007 (.322) 0.007 (.755) - -
Table 8: Stability of three simulation results measured by standard deviation. The mean values of all three simulations are reported in brackets.
Stability of the aggregative mean. The stability results of the aggregative mean are shown in Table 8. We observe a low level of variation with less than 0.7 pp for both training-based and prompting-based agents.
Dynamic stability.
model & setting t distinct-1 distinct-2 novelty-1 novelty-2
training-based $\alpha=0$ 1 0.001 (.125) 0.003 (.684) 0.001 (.036) 0.003 (.129)
2 0.001 (.120) 0.002 (.666) 0.002 (.032) 0.001 (.128)
3 0.005 (.118) 0.006 (.660) 0.001 (.034) 0.004 (.136)
4 0.001 (.116) 0.005 (.647) 0.005 (.034) 0.005 (.151)
$\alpha=2$ 1 0.002 (.147) 0.005 (.699) 0.003 (.065) 0.001 (.183)
2 0.001 (.162) 0.002 (.708) 0.005 (.081) 0.005 (.217)
3 0.007 (.175) 0.008 (.719) 0.007 (.106) 0.010 (.255)
4 0.001 (.172) 0.003 (.708) 0.005 (.126) 0.009 (.298)
prompting-based GPT-3.5 chain-prompting 1 0.019 (.338) 0.006 (.792) - -
2 0.011 (.328) 0.002 (.763) - -
3 0.017 (.317) 0.007 (.740) - -
4 0.008 (.304) 0.004 (.724) - -
Table 9: Dynamic stability of three simulation results measured by standard deviation. The highest value in each experimental setting is highlighted in bold. The mean values of all three simulations are reported in brackets.
The stability results for our dynamic statistics are shown in Table 9. We highlight the highest value in each setting in bold. For training-based agents, the results show a low variation with a maximum of 1 pp. Results at iterations 3 and 4 show a slightly higher variation for all four measures than the results at $t=1,2$. In contrast, for prompting-based agents, we observe a greater level of variation with the highest standard deviation to the level of 1.9 pp. Specifically, the results for distinct-1 exhibit more instability than other measures. This may be caused by the more diverse lexicons from prompting-based agents.
Overall, our simulations indicate high stability over the statistics, especially for training-based agents. The prompting-based agents are slightly more unstable (with a variation up to 1.9 pp) in comparison to training-based agents (with a maximum variation of 1 pp and 80% of the variation under 0.5 pp).
6.2 The effect of different learning strategies and heterogeneous models for prompting-based agents
model strategy distinct-1 distinct-2
GPT-4 positive 0.286 0.598
GPT-4 negative 0.313 0.653
GPT-4 joint (positive + negatives) 0.336 0.817
Table 10: Diversity results in aggregative mean for prompting-based agents under different learning strategies. Distinct-1 and distinct-2 are the percentage of distinct uni-/bi-grams.
Non-cooperative environments boost diversity. To examine the effect of learning strategies for prompting-based agents, we utilize the same experimental parameters as in Section 4 and conduct generation under positive-only, negative-only, and joint learning strategies (joint-prompting), respectively. As shown in Table 10, the joint learning strategy, which employs both positive and negative steps, is the most effective in terms of the diversity of the generated poetry, yielding a 5.0 pp increase in distinct-1 and over a 20 pp increase in distinct-2 compared to the positive-only strategy. Moreover, the negative-only strategy enhances diversity compared to the positive-only strategy, but to a lesser extent than the joint approach.
model distinct-1 distinct-2
GPT-4 0.336 0.817
GPT-4 + GPT-3.5 0.413 0.814
GPT-4 + GPT-3.5 + LlaMa3-7b 0.511 0.887
Table 11: Diversity results in aggregative mean for prompting-based agents with heterogeneous models.
Heterogeneous models can boost the diversity of the system. To test the effect of using non-homogeneous agents, we use a combination of various models to conduct experiments using joint-prompting defined in Section 4. As shown in Table 11, when GPT-4 is combined with GPT-3.5, the distinct-1 score increases by 7.7 pp to 0.413, while distinct-2 slightly decreases by 0.3 pp to 0.814. Incorporating LlaMa3-7b along with GPT-4 and GPT-3.5 further enhances the diversity, with distinct-1 increasing by an additional 9.8 pp to 0.511, and distinct-2 increasing by 7.3 pp to 0.887. This demonstrates the potential benefits of employing a more diverse ensemble of models.
6.3 Can different initializations lead to group-based behaviors for prompting-based agents?
As discussed in Section 5, the framework built with prompting-based agents does not exhibit any group-based behavior as expected. Considering that we initialize the agents with random samples drawn from the QuaTrain corpus, we suspect this may cause a high resemblance among the initialized poems. To examine whether an initialization with poems of more contrastive forms can produce group-based behavior, we conduct an experiment using GPT-3.5 under chain-prompting strategy where we initialize group A with poems written by Edgar Allan Poe and group B with poems written by school children under 12 years old Hipson and Mohammad (2020). An example poem from Edgar Allan Poe is “From the lightning in the sky, As it passed me flying by, From the thunder and the storm, And the cloud that took the form.” and an example poem from a school child is “Roses are red, violets are blue. I love the zoo. do you?” We implement the same process and compute the statistics. Slightly surprisingly, we observe a very similar trend for both diversity and semantic divergence to that of random initialized prompting-based agents as shown in Section 5. In terms of diversity, we notice an increase of 2 pp at iteration $t=1$ and then a decreasing trend for both distinct-1 and distinct-2. Qualitatively, at iteration $t=1$, we obtain poems from group B such as “As the sun rose, a butterfly landed softly on my hand, whispering secrets of the garden with each flutter of its delicate wings.”, which resembles the tone of a child and the imagery of a child playing in the garden. However, as $t>1$, we yield similar homogeneous poems to the case in Table 7. An example poem from group B at $t=4$ is “Beneath the starlit sky, a solitary figure stands, A soft whisper of wind caresses the quiet lands. Burdened with untold sorrows in the night so still, ‘I am but a fleeting shadow, lost in time’s skill.’ ”. The results again suggest that GPT-3.5 (also GPT-4) tends to ignore the prompts (i.e., in our case their personas) and rely more on its pretraining knowledge. This observation coincides with the work from Chuang et al. (2023); Tirumala et al. (2022) that larger models suffer more from memorization.
7 Concluding remarks
In this paper, we introduce an LLM-based multi-agent framework that incorporates not only cooperative interaction but also non-cooperative environments. We experiment with $M=4$ training-based agents trained on GPT-2 and prompting-based agents employing GPT-3.5 and GPT-4. Our evaluation with 96k generated poems shows that for training-based agents: 1) non-cooperative environments encourage diversity and novelty over iteration measured by distinct and novel n-grams; 2) training-based agents demonstrate group divergence in terms of lexicons, styles and semantics in accordance to the predefined group affiliation. Our results also indicate that for prompting-based agents: 3) the generated poetry contains very few grammatical errors with a more diverse lexicon; 4) the prompting-based framework benefits from non-cooperative environments and heterogeneous model in terms of aggregated diversity; 5) dynamically, prompting-based framework barely improves lexical diversity after the first iteration and unlike training-based agents, prompting-based agents do not show group-based divergence as expected; 6) prompting-based agents are prone to generating poetry of more homogeneous styles over time, presumably suggesting the memorization problems of LLMs.
Nowadays, more researchers have raised concerns that the use of LLMs may lead to homogeneity and uniformity of human language and knowledge Kuteeva and Andersson (2024). Empirical evidence also suggests that LLMs under the current training paradigm such as RLHF (i.e., reinforcement learning from human feedback) produce less diverse text Kirk et al. (2023); Chen et al. (2024). In this context, we believe a training paradigm shift towards a more human-like machine-learning process, particularly for creative tasks such as poetry generation, is thus necessary and meaningful. As suggested by our experiments, a more human-like (network-structured) social learning process that emphasizes non-cooperative interaction can bring in more diversity and novelty. Our results also show promise for mitigating the issues of data degeneration caused by the ‘self-consuming’ loop during modeling Wang et al. (2022a).
Future work can improve on several points. For training-based agents, enhancing inference efficiency using techniques such as speculative sampling would benefit the scaling of the framework Dekoninck et al. (2023) and thus boost diversity and novelty to a greater level. For prompting-based agents, involving more complex reasoning methods such as tree-of-thought Yao et al. (2024) into the prompting might be helpful. Extending the current framework to include an interactive combination of both training-based and prompting-based agents might be interesting to explore in which a diverse network of LLMs might bring additional generation diversity to the system.